OpenStudy (noahbred):

help please If it took 80 N-m of work to move an object 4 m, how much does the object weigh (in newtons)? 200 N 20 N 320 N .25 N

OpenStudy (12wishs):

it would be 320 newtons

OpenStudy (solomonzelman):

Hooke's Law: \(\color{#000000 }{ \displaystyle {\rm F}=kx }\) \(\color{#000000 }{ \displaystyle {\rm F}=80\color{grey}{\rm \left[N / m\right]}\cdot 4 \color{grey}{\rm \left[ m\right]}=320\color{grey}{\rm \left[N\right]}}\)

OpenStudy (solomonzelman):

Indeed, that was right!